OpenStudy (anonymous):

Two positive charges of 6.0 X 10 (exponent of -5) C are separated by 0.50 m. What force exists between the charges ??

OpenStudy (anonymous):

use F =kqq/r^2 r=.50m

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

can u please explain further how u solve it ?

OpenStudy (prakharluv):

plz check the magnitude of charge is correct or not...

OpenStudy (anonymous):

yes its correct

OpenStudy (prakharluv):

so on putting the given data on the equation \[F=k \frac{ q _{1}*q_{2} }{ r ^{2} }\] we can find out the force F=1296*10^-1N

OpenStudy (anonymous):

q1= 6.0 x 10^-5 C, q2= 6.0 x 10^-5 C distance= r= 0.50m Force= (k*q1*q2)/r^2 put the values and get the answer..

OpenStudy (anonymous):

gj with those calculators everyone!

OpenStudy (anonymous):

k is constant and value of k is 9 x 10^9 Nm^2/C^2

OpenStudy (anonymous):

okay so is that the answer - F= 1296*10 ^-1 N ? so the u wod sole it by doing F = 6.0 x 10( exponent -5) divided 0.50?

OpenStudy (anonymous):

give up.